Assume you used Analysis of Variance to compare the student debt for 100 college graduates--25 from LSU, 25 from LCC, 25 from Stetson, and 25 from University of Phoenix.

How many degrees of freedom are there for the between-group variance, the numerator of the F-statistic?

Assume you used Analysis of Variance to compare the student debt for 100 college graduates--25 from LSU, 25 from LCC, 25 from Stetson, and 25 from University of Phoenix.

How many degrees of freedom are there for the within-group variance, the denominator of the F-statistic?

Assume you used Analysis of Variance to compare the student debt for 100 college graduates--25 from LSU, 25 from LCC, 25 from Stetson, and 25 from University of Phoenix.

If the between-group variance equals 500 and the within group variance equals 50, what is the value of F
Answer #1

Since total number of observation = 100

Number of factor = k = 4

So degree of freedom of numerator

= k-1 = 4-1 = 3

The degree of freedom of denominator

= n-k = 100-4 =96

Since MST = 500

MSE = 50

F = MST/MSE = 500/50 = 10